OpenStudy (anonymous):

A triangle with a base of 6 cm and the area of 15 cm2 has a heigh of?

OpenStudy (anonymous):

area of triangle is \(\frac{1}{2}\times \text{height} \times \text{base}\)

OpenStudy (saifoo.khan):

Area = 0.5 * b * h 15 = 0.5 * 6 * h Solve for h.

OpenStudy (anonymous):

thankyou

OpenStudy (saifoo.khan):

np

OpenStudy (anonymous):

im not sure if im correct but it got 39 cm for the heigh